OpenStudy (anonymous):

At the start of the month, a brand new arcade game's play counter read 2,197. At the end of the month, it read 7,306. If the game was played for 2 1/2 minutes each time, then what was the approximate total number of minutes played on the new game for the month?

OpenStudy (lgbasallote):

the brand new arcade is second-hand?

OpenStudy (anonymous):

We begin: \[ (7306-2197)\cdot\frac{5}{2}=12772.5\text{ mins} \] And, yes, apparently so @lgbasallote

OpenStudy (lgbasallote):

that's a rip-off... :|

OpenStudy (anonymous):

Agreed...